4NH3+5O2 yields 4NO+6H2O. In a certain condition this reaction occurs and yields 45.64 NO. How many grams of NH3 and O2 are needed to make 19.6 NO.
1
Expert's answer
2017-04-20T06:46:38-0400

4NH3+5O2 yields 4NO+6H2O. In a certain condition this reaction occurs and yields 45.64 g NO. How many grams of NH3 and O2 are needed to make 19.6 g NO.

Answer:
According to the reaction, n (NH3) = n (NO), n (O2) = 5/4 n (NO).
n = m/M m = n · M
M (NO) = 28 g/mol
M (NH3) = 17 g/mol
M (O2) = 32 g/mol

n (NH3) = 19.6/28 = 0.7 mol
n (NH3) = n (NO) = 0.7 mol
m (NH3) = 0.7 · 17 = 11.9 g

n (O2) = 0.7 · 5/4 = 0.875 g
m (O2) = 0.875 · 32 = 28 g